Each side of a square is 4 m long. When each side is increased by x m, the area is doubled. Find the value of x. Thank you!

Answer by unlockmath(1688)

Hello,
We know the area is 14 Sq meters, right?
so we can set up this equation:
(x+4)^2=28
rewritten as:
x^2+8x+16=28
Let's complete the square. Subtract 16 to get:
x^2+8x=12
Add 16 to both sides:
x^2+8x+16=28
Factor:
(x+4)^2=28
Sq RT both sides:
x+4=+-sq rt 28
Subtract 4 to get:
x=-4+- sq rt 28
We get:
x=1.291 approx
We can check this out by adding 1.291 to 4 = 5.291
Square that and we get:
27.99 Approx or twice the area of the original.
Make sense?
RJ
www.math-unlock.com